# Energy-Efficient Channel Decoding for Wireless Federated Learning: Convergence Analysis and Adaptive Design

### v1

@@ III-A Effect of BER on the Convergence Rate of Wireless FL
When the learning rate is set as $\eta=\frac{1}{L\sqrt{RE}}=\frac{1}{L\sqrt{T}}$, the average expected squared norm of global stochastic gradient is bounded as follows:
$\begin{split}\frac{1}{T}\sum_{r=1}^{R}\sum_{e=1}^{E}&\mathbb{E}\|\nabla f(\bar{\mathbf{w}}_{r,e})\|^{2}\\ \leq&\frac{C_{0}}{\sqrt{T}}+\frac{C_{1}\sum_{r=1}^{R}b_{r}}{\sqrt{T}}+\frac{C_{2}}{\sqrt{T}}+\frac{C_{3}}{\sqrt{T}},\end{split}$ (6)
where $\bar{\mathbf{w}}_{r,e}\triangleq\frac{1}{n}\sum_{k=1}^{n}\mathbf{w}_{r,e}^{k}$, $C_{0}\triangleq 2L(f(\mathbf{w}_{0})-f^{*})$, $C_{1}\triangleq\frac{8L^{2}dM^{2}}{3}$, $C_{2}=\sigma_{L}^{2}(\frac{1}{n}+\frac{4(n-s)}{s(n-1)})$, $C_{3}\triangleq\frac{\sigma_{G}^{2}}{L}$, $f^{*}$ is the minimum global training loss, $b_{r}$ is the BER in the $r$-th communication round, and $\mathbf{w}_{0}$ is the initial global model.
@@
The key idea is to first quantify the BER’s impact on the model error and then introduce the model error to the convergence analysis of vanilla learning process. The detailed proof is available in the Appendix. ∎
Remark 1.
It is worthy noting that the first term on the right-hand side (RHS) of (6) captures the difference between the initial and minimum global training loss. The second term indicates how bit errors in each communication round affect learning convergence. The third term comes from the random mini-batch sampling and partial client participation, and the fourth term shows the impact of imbalanced data distribution.
According to Theorem 1, in order to guarantee the convergence rate of $\mathcal{O}(1/\sqrt{T})$, it is crucial to control the BER in each communication round. Although the explicit relation between $b_{r}$ and $r$ is difficult to obtain, we know from the above discussion that, $b_{r}$ should be a decreasing function with $r$ as otherwise the second term on the RHS of (6) cannot be diminshed with $T$ increases. For ease of implementation, we propose a simple BER regulation with $b_{r}\propto\frac{1}{r^{2}}$, which is shown in the following corollary as a sufficient condition for a practical wireless FL system to achieve the same convergence rate as that with error-free communication.
Corollary 1.
If the BER in each communication round is regulated as follows:
$b_{r}\propto\frac{1}{r^{2}},r=1,2,\cdots,R,$ (7)
the RHS of (6) diminishes to zero at the rate of $\mathcal{O}(1/\sqrt{T})$, i.e., wireless FL converges to a local optimal solution of the global training loss funcation at a rate of $\mathcal{O}(1/\sqrt{T})$.
Proof.
To prove the RHS of (6) diminishes to zero at the rate of $\mathcal{O}(1/\sqrt{T})$, it is sufficient to verify
$\lim_{T\rightarrow\infty}\frac{\sum_{r=1}^{R}b_{r}}{\sqrt{T}}=0$ (8)
when $b_{r}\propto 1/r^{2}\triangleq\delta/r^{2}$. Since
$\begin{split}0&\leq\frac{\sum_{r=1}^{R}b_{r}}{\sqrt{T}}=\delta\frac{\sum_{r=1}^{R}1/r^{2}}{\sqrt{T}}\\ &=\frac{\delta}{\sqrt{T}}\left(1+\sum_{r=2}^{R}\frac{1}{r^{2}}\right)\\ &<\frac{\delta}{\sqrt{T}}\left(1+\sum_{r=2}^{R}\frac{1}{r(r-1)}\right)=\frac{\delta}{\sqrt{T}}\left(2-\frac{1}{R}\right)\\ &=\frac{\delta}{\sqrt{T}}\left(2-\frac{E}{T}\right),\end{split}$ (9)
we conclude (8) by letting $T\rightarrow\infty$ for both sides of (9) and appying the Sandwich Theorem. ∎
Corollary 1 indicates that gradually decreasing the BER at a rate of $\mathcal{O}(1/r^{2})$ can guarantee the convergence rate of wireless FL at a rate of $\mathcal{O}(1/\sqrt{T})$. The proposed BER scaling rule achieves an improving communication quality with communication rounds. This is consistent with the solution developed in [40], which improves the communication quality by increasing the signal-to-noise ratio (SNR) with communication round.
@@ 2 Output: The converged global model $\mathbf{w}_{R}$
3 for  Communication round $r=1,2,...,R$ do
4       for Client index $k$ in $\{1,2,...,K\}$ do
5             $b_{r}\propto(1/r^{2})$ with $b_{1},b_{R}$
@@ 6             $Q=f^{-1}(b_{r})$
14            output: $\mathbf{w}_{r}^{\prime}$
15             for local training step $e=1,...,E$ do
16                  $\mathbf{w}^{k}_{r,0}=\mathbf{w}_{r}^{\prime}$
17                   $\mathbf{w}^{k}_{r,e+1}=\mathbf{w}^{k}_{r,e}-\eta_{r}\nabla f(\mathbf{w}^{k}_{r,e})$
@@ 19            $\Delta\mathbf{w}_{r}^{k}=\mathbf{w}^{k}_{r,E}-\mathbf{w}_{r,0}^{k}$
21      $\mathbf{w}_{r+1}=\mathbf{w}_{r}+\sum_{k=1}^{K}p_{k}\Delta\mathbf{w}_{r}^{k}$
@@ III-C Decoding Iteration Control
$Q_{\color[rgb]{0,0,0}r}=\frac{1}{\chi_{0}}\mathcal{O}\left(2\log\ r\right)+\frac{1}{\chi_{0}}\mathcal{O}\left(\frac{1}{\epsilon}\log\frac{1}{\epsilon}\right),$ (13)
which provides a guideline to select the maximum number of LDPC decoding iterations in the $r$-th communication round. Specifically, $Q_{r}$ can be set as $\alpha\log r+\beta$, where $\alpha$ and $\beta$ are constants. However, (13) cannot be directly utilized because $\alpha$ and $\beta$ are code-specific and depend on the receive SNR as well as channel decoding algorithms. Thus, we propose a practical approach by a lookup table, which captures the relationship between BER and the maximum number of decoding iterations at different levels of receive SNR. Entries in the lookup table can be obtained through empirical measurements specific to the LDPC code used in the wireless FL system. Fig. 3 shows the sample simulation results for a rate 1/2 LDPC code with a block length of 1008 bits and binary phase shift keying (BPSK) modulation over an additional white Gaussian noise (AWGN) channel with $\text{SNR}=1.5$ and 2.5 dB. Data in this figure can be stored in a three-dimensional lookup table [46]. From the lookup table, we can get the corresponding maximum number of decoding iteration for a target BER for clients with different receive SNR.
@@
The wireless FL procedures with adaptive LDPC decoding are summarized in Algorithm 1. The algorithm starts with an initial target BER, denoted as $b_{0}$, which is set to a relatively high value, i.e., $10^{-1}$ in our experiments. The target BER at the last communication round, denoted as $b_{R}$, is set as a small value, i.e., $10^{-4}$ in our experiments. The concrete expression of $b_{r}$ in our experiment is $b_{r}=\frac{0.0999R^{2}}{(R^{2}-1)r^{2}}+\frac{10^{-4}R^{2}-0.1}{R^{2}-1}$. For each communication round, the BER is scaled according to (7), and the corresponding maximum number of LDPC decoding iterations in that round is obtained from the lookup table.
@@
Note that the number of entries in the lookup table depends on the possible discrete BER and SNR values. Since this number is not large, the memory overhead for storing the lookup table is expected to be small. Additionally, reading data from the lookup table takes constant time and does not introduce additional LDPC decoding latency.
@@ IV Experimental Results
In this section, we validate the effectiveness of the proposed adaptive LDPC decoding algorithm via numerical experiments. First, we introduce the experimental settings. Then, the performance of the proposed algorithm, including test accuracy and energy consumption, is compared with a fixed-iteration LDPC decoding baseline.
Figure 4: Fixed decoding schemes: On the mobile side, setting the LDPC maximum decoding iterations to 6, 12, 18, 24, and 52 can achieve a BER of $10^{-1}$, $10^{-2}$, $10^{-3}$, $10^{-4}$, and $10^{-5}$, respectively.
@@ IV-A Experimental Settings
System Setups: We consider an FL system with 10 clients. Clients are assumed active in every training round and they communicate with the server using rate 1/2 min-sum LDPC code with a code length of 1008[47]. Besides, binary phase shift keying (BPSK) is adopted as the modulation scheme. In each training round, the number of local SGD steps is set as 5 with the learning rate and batch size as 0.01 and 64, respectively.
Datasets: We conduct experiments on three dataset, including MNIST[48], Fashion-MNIST[49], and CIFAR-10[50]. The training datasets were distributed among mobile devices in an independent and identically distributed (IID) manner. We further consider one non-IID experiment on Fashion-MNIST where each client was assigned only two categories of training data, similar to the setup in [27]. The test datasets were utilized for validation on the server side.
@@ IV-B Learning Performance of Conventional Schemes
Figure 7: Experiment for LeNet-300-100 on MNIST. Compared with the conventional fixed schemes, our method that adaptively increases the maximum decoding iteration to 24 can achieve the same final accuracy of 98%, and reduce the decoding energy by 6.7mJ.
@@
Figure 8: Experiment for vanilla CNN on Fashion-MNIST. Compared with the conventional fixed schemes, our method that adaptively increases the maximum decoding iteration to 24 can achieve a similar final accuracy of 91%, and reduce the decoding energy by 14.1mJ.
In this subsection, the experimental results for conventional schemes that use fixed decoding iterations are presented. As shown in Fig. 4, the fixed decoding iterations lead to fixed BERs in the whole training process.
Next, the impact of fixed BER on FL performance was investigated. Fig. 5 summarizes the test accuracy with different fixed number of maximum decoding iterations. It can be observed that for all four FL tasks, when BER is higher than $10^{-3}$, the learning performance was significantly impacted. Lowering the BER to $10^{-4}$ leads to improved accuracy and convergence towards perfect communication levels. However, further decreasing the BER to $10^{-5}$ will not yield any significant improvement in learning performance.
IV-C Learning Performance of the Proposed Approach
In this subsection, the proposed BER scaling algorithm is compared to the fixed BER scheme. The proposed approach uses a scaling BER of $\mathcal{O}(1/r^{2})$ by controlling the maximum decoding iteration in each communication round based on a lookup table that summarizes the relationship between the BER and decoding iterations.
@@ IV-C Learning Performance of the Proposed Approach
Figure 9: Experiment for 7-layer CNN on CIFAR-10. Compared with the conventional fixed schemes, our method that adaptively increases the maximum decoding iteration to 24 can achieve a similar final accuracy of 77.6%, and reduce the decoding energy by 190mJ.
@@
Figure 10: Experiment for vanilla CNN on non-IID Fashion-MNIST. Compared with the conventional fixed schemes, our method that adaptively increases the maximum decoding iteration to 24 can achieve a similar final accuracy of 54.1%, and reduce the decoding energy by 13.9mJ.
Specifically, the learning process is assigned with an initial BER value of $10^{-1}$ and a target BER value of $10^{-4}$. Based on the BER scaling rule of $\mathcal{O}(1/r^{2})$, the desired BER at every communication round can be calculated. Then based on the lookup table, we can control the LDPC maximum decoding iteration for every communication round. In the experiments, the initial and final maximum decoding iterations are set to 6 and 24, respectively. The learning performance and energy consumption of the proposed adaptive scheme are compared to those of fixed schemes with fixed BER of $10^{-4}$. The corresponding fixed maximum decoding iterations is set to 24. To mitigate energy consumption, the LDPC decoder employs an early termination scheme, where the decoding process stops when the bitstream in a frame is deemed correct. Thus, only frames that cannot be correctly decoded require the maximum decoding iteration. Fig. 6 shows an example for the early termination scheme by using the experiment of vanilla-CNN on Fashion-MNIST, illustrating the actual decoding iterations for each bitstream. It can observed that only a few bitstreams use up the maximum decoding iterations which is set to 52. The average decoding iterations is the average number of real consumed decoding iterations. Given the decoding energy is proportional to the decoding iterations, the decoding energy consumption can be reflected by the average number of decoding iterations used for the FL task.
@@
Fig. 7 summarises the experimental results for LeNet-300-100 on MNIST. Fig. 7(a) shows how the maximum number of LDPC decoding iterations varies with the FL communication rounds for the proposed approach compared with other fixed schemes. The traditional schemes adopt a fixed number of maximum decoding iterations, resulting in a fixed BER, while the proposed scheme adopts a roughly linearly increasing number of maximum decoding iterations to achieve the designed BER. Fig. 7(b) shows the learning accuracy curves for different schemes. Setting the fixed maximum decoding iteration to 6, which corresponds to a BER of $10^{-1}$, severely affects the training accuracy. Similarly, when the maximum decoding iteration is set to 12, resulting in a BER of about $10^{-2}$, the training still fails to converge. With a maximum decoding iteration of 18, corresponding to a BER of about $10^{-3}$, the training can converge, but with a noticeable loss in accuracy. The training converges well to high learning accuracy when the maximum decoding iteration is increased to 24, resulting in a BER of $10^{-4}$. Further increasing the maximum decoding iteration to 52 does not lead to further accuracy improvement.
@@
The proposed scheme, which approximately increases the maximum decoding iteration from 6 to 24 in a linear manner, enables the training accuracy to improve rapidly and converge to the same level as the fixed scheme with 24/52 iterations. Fig. 7(c) illustrates how the average number of decoding iterations varies during training. The proposed scheme can reduce the average number of iterations and thus the overall energy consumption. Fig. 7(d) depicts the BER of the proposed scheme, which scales as $\mathcal{O}(1/r^{2})$ from $10^{-1}$ to $10^{-4}$. Fig. 7(e) shows the final accuracy of different schemes. It can be observed that the proposed approach achieves the same accuracy as the fixed schemes that always maintain a very low BER of $10^{-4}$ and $10^{-5}$ throughout the training process. Fig. 7(f) illustrates the decoding energy consumption on the mobile side for the different schemes. It is obtained by multiplying the total number of LDPC decoding iterations with the LDPC decoding energy efficiency reported in [32]. Compared to the fixed schemes using a maximum decoding iteration of 24, the proposed scaling scheme can reduce the overall decoding energy by 19.4%.
Figs. 8, 9, and 10 present the experimental results for Fashion-MNIST, CIFAR-10, and non-IID Fashion-MNIST, respectively. The results are similar to those obtained by the MNIST experiment and show that the adaptive scheme can achieve similar or even slightly better accuracy compared to the fixed schemes. The observation of slightly better learning accuracy was also reported in [52, 53, 54], which suggests that adding noise to the model could enhance its generalization ability, and annealed noise could help achieve better learning accuracy. Overall, the proposed scheme reduces the LDPC decoding energy for Fashion-MNIST, CIFAR-10, and non-IID Fashion-MNIST by 18.3%, 20.1%, and 18.0%, respectively. Table III provides a summary of the experimental data.
V Conclusion
In this paper, we considered the energy consumption issue in federated learning (FL) over digital communication networks. The study recognized that channel decoding contributes to a large portion of the energy consumption for mobile FL, which has not been previously highlighted. It was shown that, due to the robustness of FL to communication errors, properly control the BER over different communication rounds can achieve the same convergence rate as error-free FL. Based on this analysis, we proposed an energy-efficient LDPC decoding iteration control algorithm. Experimental results demonstrated that the proposed method can reduce approximately 20% the total energy consumed by LDPC decoding, while maintaining the same or even a slightly higher learning accuracy compared to the scheme without decoding iteration control.

### v2

@@ III-A Effect of BER on the Convergence Rate of Wireless FL
With Assumptions 1-5 and a sufficient large value of $R$ such that $\eta=\frac{1}{L\sqrt{RE}}$ satisfies (8), (9) can be simplified as follows:
$\begin{split}\frac{1}{T}&\sum_{r=0}^{R-1}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\\ \leq&\frac{C_{0}}{\sqrt{T}}+\frac{C_{1}\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}}{\sqrt{T}}\\ &+\frac{C_{2}\sum_{r=0}^{R-1}(b_{r}(1-b_{r})^{N-1})^{2}}{\sqrt{T}}+\frac{C_{3}}{\sqrt{T}}+\frac{C_{4}}{T},\end{split}$ (10)
where $C_{0}\triangleq 2L(\mathbb{E}f(\mathbf{w}_{0})-f^{*})$, $C_{1}\triangleq\frac{L^{2}(L+2)dM^{2}(4^{N}-1)}{3(2^{N}-1)^{2}}$, $C_{2}\triangleq LdM^{2}$, $C_{3}\triangleq\frac{(L+1)(\sigma_{L}^{2}+\sigma_{G}^{2})}{n}$, $C_{4}\triangleq\frac{L^{2}\sigma_{L}^{2}(n+1)(E-1)}{n}$.
According to Theorem 1, in order to guarantee the convergence rate of $\mathcal{O}(1/\sqrt{T})$, it is crucial to control the BER in each communication round such that the RHS of (10) reduces to zero at a rate of $\mathcal{O}(1/\sqrt{T})$. Since the explicit relationship between $b_{r}$ and $r$ is difficult to obtain, we propose a simple BER regulation method in Corollary 1. As such, wireless FL systems can achieve the same convergence rate as those with error-free communication.
Corollary 1.
If the BER in each communication round is regulated as
$b_{r}\propto\frac{1}{(r+1)^{2}},r=0,\cdots,R-1,$ (11)
the RHS of (10) diminishes to zero at a rate of $\mathcal{O}(1/\sqrt{T})$, i.e., wireless FL converges to a local optimal solution of the global training loss funcation at a rate of $\mathcal{O}(1/\sqrt{T})$.
@@
When $b_{r}\propto\frac{1}{(r+1)^{2}}\triangleq\frac{\vartheta}{(r+1)^{2}}$, to prove the RHS of (10) diminishes to zero at a rate of $\mathcal{O}(1/\sqrt{T})$, it is sufficient to verify
$\lim_{T\rightarrow\infty}\frac{\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}}{\sqrt{T}}=0$ (12)
due to the fact that $0\leq\left(b_{r}(1-b_{r})^{N-1}\right)^{2}\leq b_{r}(1-b_{r})^{N-1}\leq 1$ for $0\leq b_{r}\leq 1$. Specifically, since
$\begin{split}0&\leq\frac{\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}}{\sqrt{T}}\leq\frac{\sum_{r=0}^{R-1}b_{r}}{\sqrt{T}}=\vartheta\frac{\sum_{r=0}^{R-1}1/(r+1)^{2}}{\sqrt{T}}\\ &=\frac{\vartheta}{\sqrt{T}}\left(1+\sum_{r=1}^{R-1}\frac{1}{(r+1)^{2}}\right)\\ &<\frac{\vartheta}{\sqrt{T}}\left(1+\sum_{r=1}^{R-1}\frac{1}{r(r+1)}\right)=\frac{\vartheta}{\sqrt{T}}\left(2-\frac{1}{R}\right)\\ &=\frac{\vartheta}{\sqrt{T}}\left(2-\frac{E}{T}\right),\end{split}$ (13)
we conclude the proof by letting $T\rightarrow\infty$ for both sides of (13) and applying the Sandwich Theorem. ∎
Corollary 1 indicates that gradually decreasing the BER at a rate of $\mathcal{O}(1/(r+1)^{2})$ can guarantee wireless FL converges at a rate of $\mathcal{O}(1/\sqrt{T})$. The proposed BER regulation rule achieves an improving communication quality with communication rounds. This is consistent with the solution developed in [36], which improves the communication quality by increasing the signal-to-noise ratio (SNR) with communication round.
@@ 2 Output: The converged global model $\mathbf{w}_{R}$
3 for  Communication round $r=0,\cdots,R-1$ do
4       for Client index $k=1,\cdots,n$ do
5             $b_{r}\propto(1/(r+1)^{2})$ with $b_{0},b_{R-1}$
@@ 6             $Q=f^{-1}(b_{r})$
14            output: $\mathbf{\tilde{w}}_{r}$
15             for local training step $e=0,...,E-1$ do
16                  $\mathbf{w}^{k}_{r,0}=\mathbf{\tilde{w}}_{r}$
17                   $\mathbf{w}^{k}_{r,e+1}=\mathbf{w}^{k}_{r,e}-\eta\nabla f_{k}(\mathbf{w}^{k}_{r,e};\xi^{k}_{r,e})$
@@ 19            $\Delta\mathbf{w}_{r}^{k}=\mathbf{w}^{k}_{r,E}-\mathbf{w}_{r,0}^{k}$
21      $\mathbf{w}_{r+1}=\mathbf{w}_{r}+\frac{1}{n}\sum^{n}_{k=1}\Delta\mathbf{w}_{r}^{k}$
@@ III-C Decoding Iteration Control
$Q_{r}=\frac{1}{\chi_{0}}\mathcal{O}\left(2\log\ r\right)+\frac{1}{\chi_{0}}\mathcal{O}\left(\frac{1}{\epsilon}\log\frac{1}{\epsilon}\right),$ (17)
which provides a guideline to select the maximum number of LDPC decoding iterations in the $r$-th communication round. Specifically, $Q_{r}$ can be set as $\alpha\log r+\beta$, where $\alpha$ and $\beta$ are constants. However, (17) cannot be directly utilized because $\alpha$ and $\beta$ are code-specific and depend on the receive SNR as well as channel decoding algorithms. Thus, we propose a practical approach by the mapping relationship which captures the relationship between BER and the maximum number of decoding iterations at different levels of receive SNR. Entries in the mapping relationship can be obtained through empirical measurements specific to the LDPC code used in the wireless FL system. Fig. 5 shows the exampled simulation results for a rate 1/2 LDPC code with a block length of 1008 bits and binary phase shift keying (BPSK) modulation over an additional white Gaussian noise (AWGN) channel with $\text{SNR}=1.5$ and 2.5 dB. Data in this figure can be stored in memory [43]. From the mapping relationship, we can get the corresponding maximum number of decoding iteration for a target BER for clients with different receive SNR.
@@
The wireless FL procedures with adaptive LDPC decoding are summarized in Algorithm 1. The algorithm starts with an initial target BER, denoted as $b_{0}$, which is set to a relatively high value, i.e., $10^{-1}$ in our experiments. The target BER at the last communication round, denoted as $b_{R-1}$, is set as a small value, i.e., $10^{-4}$ in our experiments. The concrete expression of $b_{r}$ in our experiment is $b_{r}=\frac{0.0999R^{2}}{(R^{2}-1)(r+1)^{2}}+\frac{10^{-4}R^{2}-0.1}{R^{2}-1}$. For each communication round, the BER is scaled according to (11), and the corresponding maximum number of LDPC decoding iterations in that round is obtained from the mapping relationship.
@@
Note that the number of entries in the mapping relationship depends on the possible discrete BER and SNR values. Since this number is not large, the memory overhead for storing the mapping relationship is expected to be small. Additionally, reading data from the memory takes constant time and does not introduce additional LDPC decoding latency.
@@ IV Experimental Results
In this section, we validate the effectiveness of the proposed adaptive LDPC decoding algorithm via numerical experiments. First, we introduce the experimental settings. Then, the performance of the proposed algorithm, including test accuracy and energy consumption, is compared with a baseline with a fixed number of LDPC decoding iterations.
Figure 6: Fixed decoding schemes: On the client side, setting the LDPC maximum decoding iterations to 6, 12, 18, 24, and 52 can achieve a BER of around $10^{-1}$, $10^{-2}$, $10^{-3}$, $10^{-4}$, and $10^{-5}$, respectively.
@@ IV-A Experimental Settings
System Setups: We consider an FL system with 10 clients. All clients are assumed to be selected in every communication round and they communicate with the server using rate 1/2 min-sum LDPC code with a code length of 1008 bits[44]. Besides, binary phase shift keying (BPSK) is adopted as the modulation scheme. In each communication round, the number of local SGD steps is set as 5 with the learning rate and batch size as 0.01 and 64, respectively. The receive SNR of all the wireless channels is assumed to be 2.5 dB.
Datasets: We conduct experiments on three dataset, including MNIST[45], Fashion-MNIST[46], and CIFAR-10[47]. The training datasets are distributed among clients in an IID manner. We further consider one experiment where the Fashion-MNIST data are non-IID among clients. In particular, each client was assigned with two categories of training data, similar to the setup in [24]. The test datasets are used for validation at the server.
@@ IV-B Learning Performance of Conventional Schemes
Figure 9: Experiment for LeNet-300-100 on MNIST dataset. Compared with the conventional fixed schemes, our method that adaptively increases the maximum decoding iteration to 24 can achieve the same final accuracy of 98%, and reduce the decoding energy by 6.7mJ.
@@
Figure 10: Experiment for vanilla CNN on IID Fashion-MNIST dataset. Compared with the conventional fixed schemes, our method that adaptively increases the maximum decoding iteration to 24 can achieve a similar final accuracy of 91%, and reduce the decoding energy by 14.1mJ.
In this subsection, experimental results for the baseline scheme are presented. As shown in Fig. 6, it leads to approximately constant BERs over the training process.
Next, we investigate the impact of a constant BER on the performance of wireless FL. Fig. 7 summarizes the test accuracy with different maximum numbers of decoding iterations. It can be observed that for all the four FL tasks, when the BER is higher than $10^{-3}$, the learning performance was significantly impaired. Lowering the BER to $10^{-4}$ leads to improved accuracy and faster convergence towards wireless FL with error-free communication. However, further decreasing the BER to $10^{-5}$ does not yield any significant improvement in test accuracy.
IV-C Learning Performance of the Proposed Scheme
@@
Figure 11: Experiment for 7-layer CNN on CIFAR-10 dataset. Compared with the conventional fixed schemes, our method that adaptively increases the maximum decoding iteration to 24 can achieve a similar final accuracy of 77.6%, and reduce the decoding energy by 190mJ.
@@
Figure 12: Experiment for vanilla CNN on non-IID Fashion-MNIST dataset. Compared with the conventional fixed schemes, our method that adaptively increases the maximum decoding iteration to 24 can achieve a similar final accuracy of 54.1%, and reduce the decoding energy by 13.9mJ.
In this subsection, the proposed adaptive LDPC decoding scheme is compared to the baseline scheme. Specifically, the learning process of the proposed scheme is assigned with an initial BER of $10^{-1}$ and a final BER value of $10^{-4}$. The desired BER at every communication round is calculated via (11) and the maximum number of LDPC decoding iterations is determined according to the mapping with BER as stated in Section III. In the experiments, the initial and final maximum numbers of LDPC decoding iterations are thus set to 6 and 24, respectively. For the baseline scheme, a maximum of 24 LDPC decoding iterations are used in each communication round, which corresponds to a BER of $10^{-4}$.
Fig. 8(a) shows an example on the Fashion-MNIST (IID) datasets to illustrate the actual number of decoding iterations for each frame of data bits. It can observed that only a few frames exhaust all the decoding iterations which is set to a large value of 52. Note that the decoding energy consumption is proportional to the actual number of decoding iterations used for the FL tasks. From the distribution on the actual number of decoding iterations shown in Fig. 8(b), it is observed that 18 decoding iterations are sufficient for most cases, implying the significant reduction of energy consumption.
@@
Fig. 9 summarizes the experimental results for LeNet-300-100 on the MNIST (IID) dataset. Specifically, Fig. 9(a) shows how the maximum number of LDPC decoding iterations varies with communication rounds in the proposed scheme. While the baseline scheme with a fixed number of maximum decoding iterations results in a constant BER, the proposed scheme achieves the desired BER by adopting a linearly increasing maximum number of decoding iterations. Fig. 9(b) shows the test accuracy achieved by different schemes. It is observed that setting the maximum number of decoding iterations to 6 (which corresponds to a BER of $\sim 10^{-1}$) severely affects the learning accuracy. Similarly, when the maximum number of decoding iterations is set to 12 (which corresponds to a BER of $\sim 10^{-2}$), the training process still fails to converge. By setting the maximum number of decoding iterations as 18 (which corresponds to a BER of $\sim 10^{-3}$), the training process converges, but with a noticeable loss in test accuracy compared to the case when the maximum number of decoding iterations is 24 (which corresponds to a BER of $\sim 10^{-4}$). However, further increasing the maximum number of decoding iterations to 52 does not lead to much accuracy improvement.
@@
The proposed scheme, which linearly increases the maximum number of decoding iterations from 6 to 24, achieves the same test accuracy as those with fixed 24 or 52 decoding iterations. Fig. 9(c) illustrates how the average number of decoding iterations over all DNN parameters varies during the training process. The proposed scheme can reduce the average number of decoding iterations and thus the overall energy consumption. Fig. 9(d) depicts the BER of the proposed scheme over communication rounds, which scales as $\mathcal{O}(1/(r+1)^{2})$ from $10^{-1}$ to $10^{-4}$. Fig. 9(e) shows the final test accuracy of different schemes at the 50-th communication round. It can be observed that the proposed scheme achieves the same accuracy as the baseline schemes that always maintain a very low BER of $10^{-4}$ and $10^{-5}$ throughout the training process. Fig. 9(f) illustrates the decoding energy consumption on clients under different schemes. It is obtained by multiplying the total number of LDPC decoding iterations with the LDPC decoding energy efficiency reported in [29]. Compared to the baseline scheme with a maximum of 24 decoding iterations, the proposed scheme reduces the overall decoding energy consumption by 19.4%.
@@
Figs. 10, 11, and 12 present the experimental results for Fashion-MNIST (IID), CIFAR-10 (IID), and Fashion-MNIST (non-IID) datasets, respectively. The results are similar to those on the MNIST (IID) dataset in Fig. 9, showing that adapting the maximum number of LDPC decoding iterations can achieve similar or even slightly better accuracy compared to the baseline scheme. This observation was also reported in [49, 50, 51], which suggests that adding annealed noise to the DNN model could enhance its generalization ability. Moreover, the proposed scheme reduces the LDPC decoding energy consumption on the Fashion-MNIST (IID), CIFAR-10 (IID), and Fashion-MNIST (non-IID) datasets by 18.3%, 20.1%, and 18.0%, respectively. Table III provides a summary of the experimental data.
V Conclusion
In this paper, we considered the energy consumption issue in wireless federated learning (FL) over digital communication networks. Our study recognized that channel decoding contributes to a large portion of the energy consumption at mobile clients in wireless FL systems. To achieve energy-efficient wireless FL, we first show that by properly controlling the bit error rate (BER) over different communication rounds, wireless FL can achieve the same convergence rate as FL with error-free communication. Based on this analysis, we proposed an energy-efficient LDPC decoding scheme to regulate the maximum number of decoding iterations over communication rounds. Experimental results demonstrated that the proposed method can reduce approximately 20% of the total energy consumed by LDPC decoding, while maintaining the same or even a slightly higher learning accuracy compared to the scheme without adaptive decoding iteration control.